[figure: 9.7 cm]

ORBITAL ORDERING AND FLUCTUATIONS IN TRANSITION METAL OXIDES

Andrzej M. Oleś

Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

A short review of the frustrated magnetic interactions in transition metal oxides with orbital degrees of freedom is presented. The classical phase diagrams of e_{g} systems show a competition between different phases, with spin, orbital and mixed spin-and-orbital excitations, and strongly enhanced quantum fluctuations [1]. The orbital order explains a quasi-one-dimensional antiferromagnetic (AF) interactions of spins S = 1/2 in KCuF_{3}, and the measured superexchange between spins S = 2 in the A-AF phase of LaMnO_{3}. The spectral functions of a single hole in LaMnO_{3} depend strongly on the type of the underlying orbital order [2]. Macroscopic hole doping x ≃ 0.3 promotes an orbital liquid in the ferromagnetic manganites, and the observed increase of the spin stiffness with x is well explained by e_{g} electron correlations [3]. Quantum fluctuations of t_{2g} orbitals play a prominent role in early transition metal oxides — they explain the isotropic AF order in LaTiO_{3}, and enhance ferromagnetic interactions in the C-AF phase of cubic vanadates. Finally, the t_{2g} orbital fluctuations stand behind the exotic magnetic properties of YVO_{3}, and stabilize the C-AF over G-AF order at finite temperatures [4].

[1] L.F. Feiner, A.M. Oleś, and J. Zaanen, Phys. Rev. Lett. 78, 2799 (1997).

[2] J. Bała et al., Phys. Rev. Lett. 87, 067204 (2001).

[3] A.M. Oleś and L.F. Feiner, Phys. Rev. B 65, 052414 (2002).

[4] G. Khaliullin, P. Horsch, and A.M. Oleś, Phys. Rev. Lett. 86, 3879 (2001).

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : invited talk

Corresponding author : Andrzej M. Oleś

Address for correspondence : Institute of Physics Jagellonian University Reymonta 4 PL-30059 Kraków

Email address : amoles@if.uj.edu.pl